OpenStudy (anonymous):

how do you do -6 - 2k + 11=18k

OpenStudy (anonymous):

Well, first you have to do what's called 'combining like terms'. All constants, without 'k', are 'like terms'. They can be added together.

OpenStudy (anonymous):

Similarly, the terms that have a 'k' can be added together.

OpenStudy (anonymous):

So first, you combine the constants on the left -- what do you get?

OpenStudy (anonymous):

Do you do 2k + 18k ?

OpenStudy (anonymous):

yes, if you add 2k to both sides, then the right has 2k plus 18k. You also combine the non-k terms, called constant terms, on the left. What do you end up with then?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

exactly! So you have 5 on one side, and 20k on the other, right? Then divide both sides by 20, what do you get?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Almost -- you're dividing 20 by 5, but you actually want to do the opposite. You have 20k = 5 So if you divide both sides by 20, you get 20k/20 = 5/20 k = 5/20 k = 1/4

OpenStudy (anonymous):

thanks !